HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

NEW SUPERSONIC JET NEAR IR SPECTRA OF AMMONIA AND CHLOROMETHANE

\underline{O. VOTAVA}, K. BURIAN, and J. RAKOVSKÝ, J. Heyrovský Institute of Physical Chemistry of the CAS, Dolejškova 2155/3, Prague 8, 182 23, Czech Republic

We report new high resolution near IR spectra of chloromethane (CH_{3}Cl) and ammonia (NH_{3}): Specifically for chloromethane 2ν_{4}^{0} (6007 cm^{−1} - 6033 cm^{−1}) and 2ν_{1} (5850 cm^{−1} - 5910 cm^{−1}) vibratioal bands have been recorded and two combination bands of ammonia were recorded in ranges 5990 cm^{−1} - 6090 cm^{−1} and 7590 cm^{−1} - 7710 cm^{−1} All data were obtained with tunable diode laser spectrometer coupled to pulsed, slit nozzle (0.1mm x 40mm) supersonic jet. Observed absorption spectra exhibit low rotational and Doppler temperature down to 20 K and as a result are significantly less congested compared to room temperature data. Spectral line positions have been determined with accuracy of 3x10^{−4}cm^{−1} by referencing to methane transitions in the same spectral regions.

Ammonia spectra are compared with existing FTIR data for assignment verifications and also with ExoMol CoYuTe linelist ^{1}. Agreement with the CoYuTe prediction is rather exeptional, with standart devidation between measured and predicted line position 0.056 cm^{−1} in the 7600 cm^{−1} band. The chloromethane spectra are partially assigned using effective Hamiltonnian approach and verified with GSCD. The data are also compared with theoretical Exomol linelists ^{2}. The overall agreement between the theory and experiment is lower than in the case of ammonia: The band origin for the CH_{3} ^{35}Cl ν_{1} + ν_{4} vibration is about 1.64 cm^{−1} higher than the experimental value.

1P.A. Coles et al., Monthly Notices of the Royal Astronomical Society\underline{490}, 4638 (2019).

2A. Owens et al., Monthly Notices of the Royal Astronomical Society \underline{479}, 3002 (2018).

p-number: p241  Submitted on Thu, 22 Jun 2023 16:23:17 +0200